OpenStudy (anonymous):

Solve. The formula S = P - 0.2P gives the sale price, S, of a shirt that was marked down 20% from the original price, P. Find the sale price of a shirt that originally cost $25.

OpenStudy (anonymous):

S=P-0.2P P=25 S=25-0.2(25) S=20